im brandon, i need help with substitution. problem is x-2y+2=0 and 2x-3y-2=0. i dont understand this one. please help?
Answer by Alan3354(69443)

x-2y+2=0
-2x-3y-2=0
---------
Solve for x in the 1st eqn
x-2y+2 = 0
x = 2y-2
----
Sub that for x in the other eqn.
-2x-3y-2=0
-2(2y-2) - 3y - 2 = 0
-4y+4 - 3y = 2
-7y = -2
y = 2/7
----------
x = 2y-2
x = 4/7-2
x = -10/7